Answer the following five questions related to code output (Q2 i through m) What is the output of the following C+ coaer int numl; int num2; int *p = &num1 ; p = &num2 ; *p 25; num1 num2 + 6; p = &num! ; num2 = 73; *p = 47; cout << *p << << numl <<<< num2 << endl; What is the output of the following C++ code? int length; int *width; length = new int; * length = 5; width = length; length new int; * length= 2 * (* width); cout << *length <<<<width << << (*length) *(*width) << endl; What is the output of the following C++ code? int * first new int; int second; *first 85; second = first; *secondsecond + *first; first = new int; * first = * second - 100; cont << *first<< << * second << endl;Please help me solve the following. Thank you in advance :)
Answer #1

I) Initially p is pointing to num1 but then it points to num2
p changes value of num2 (as it points to it) to 25 i.e num2 = 25
num1 = 25 + 6 = 31 , num2 will be 73 and num1 . will be 37 as p is pointing to num1 now
So, final Output is 37 37 73


II)
10 5 50

length is 5 so width also points to same location, then length gets new memory location and assigns 2*5 = 10


III)*second = 160 , then *first = 60

So, final Output is 60 160


IV) Initally the array will contain
8 10 12 14 16

temp will change the list and increment each element by 3

11 13 15 17 19

So, final Output is 11 13 15 17 19


V)
IV) Initally the array myList will contain
3 4 6 9 13

YourList = 7 8 10 13 17 10 6 3 1 0


So, final Output is myList = 3 4 6 9 13
So, final Output is YourList = 7 8 10 13 17 10 6 3 1 0